# Passive iFIR filters for data-driven control

### v1

@@ IV-A Fitting of passive FIR filters
Results are shown in Figure 2. The top-left plot shows that all three approaches give perfect fitting. The top-right plot shows results for conservative $\epsilon$ (23d). Indeed, this parameter can be used to enforce excess of passivity, to compensate for shortages in the plant. Left and right bottom plots are related to the non-passive targets $C_{2}$ and $C_{3}$. All three approaches achieve a passive FIR filter. However, as expected, the matching with the target systems degrades as the shortage of passivity increases.
IV-B Velocity control of a compliant mechanical system
@@
We now develop a full iFIR controller design. Consider the two-cart plant of Figure 3, modelling a truck pulling a heavy load. The links between the truck and the load are approximated by a parallel combination of a linear spring and a linear damper. The cart masses are $m_{1}=3.0\mbox{ kg},m_{2}=0.5\mbox{ kg}$. $k_{12}=1.0$ Ns/m is the spring stiffness and $c_{12}=1.05$ Ns/m is the damping coefficient. Each cart is also affected by additional dissipation of the form $c\dot{x}_{i}$, where $c=0.5$ Ns/m and $\dot{x}_{i}$ is the absolute velocity of the cart $i\in\{1,2\}$. The block diagram for this closed-loop system is represented in Figure 1. $u$ is the applied force and $y$ is the velocity output.
@@
We consider the reference model
$M_{r}(s)$ $=\frac{1+0.1s}{(1+2\zeta sT_{1}+s^{2}T_{1}^{2})}\qquad\zeta=1,T_{1}=0.5.$
The plant and the reference model are discretized with a sampling time of $0.05$s. The plant is excited in open-loop with $u(t)=\sum_{i=1}^{10}\sin(\omega_{i}t)$ where $\omega_{i}$ linearly spans the frequency range of $[0.5,10]$ rad/s, and $t\in\{0,...,N-1\}$. The data $(e,u)$ for the controller are obtained as discussed in Section II and filtered with the filter $M_{r}(z)$.
The computation times for several iFIR controllers of order $m\in\{50,150,250,350\}$ are summarized in Table I. For $m=350$, the KYP approach takes more than one hour. This shows the advantage of (3) and (4) over (1).
@@
KYP 0.25 s 49.04 s 769.06 s ///
Toeplitz, $n=m$ 0.07 s 0.51 s 2.81 s 4.88 s
Toeplitz, $n=1.5m$ 0.12 s 1.88 s 5.21 s 12.18 s
Toeplitz, $n=2m$ 0.17 s 3.36 s 10.62 s 61.14 s
PR, $M=m$ 0.05 s 0.14 s 0.17 s 0.25 s
PR, $M=1.5m$ 0.04 s 0.14 s 0.17 s 0.24 s
PR, $M=2m$ 0.05 s 0.14 s 0.20 s 0.26 s
TABLE I: The computation times of constraints (1), (3), and (4).
@@
The Nyquist plots of the FIR part of the iFIR controllers are shown in Figure 4. We have taken $m=350$, $n=2m$ for (3) and $M=2m$ for (4). $\epsilon$ is chosen heuristically. The differences between the two iFIR controllers are negligible. On the right-part of the complex plane, both passive iFIR controllers (black) show excellent matching with the optimal unconstrained ones (blue) .
The iFIR controllers are tested in simulation, as shown in Figure 5. The figure shows a comparison between iFIR controllers and the optimal PID controller $C(z)=0.8051+\frac{4.409T_{s}}{1-z^{-1}}+0.0068\frac{z-1}{zT_{s}}$. The PID gains are tuned using the VRFT approach (2). The gains are constrained to be positive for passivity. The closed-loop response of the two iFIR controllers is superior, with an almost exact matching to the target model. This is further illustrated by the Bode diagrams of the closed loop transfer functions, in Figure 7.
@@
Figure 4: Nyquist diagrams of the FIR part of the iFIR controller of Section IV-B. Left: blue - unconstrained FIR (2); black - constrained FIR (2), (3). Right: blue - unconstrained FIR (2); black - constrained FIR (2), (4).
@@
Figure 5: Step responses. Red: target response. Black: closed-loop response with iFIR controller. Blue-continuous: closed-loop response with PID. Blue-dashed: open loop. Left plot: iFIR (2), (3). Right plot: iFIR (2), (4).
We also design and test the iFIR controller on a nonlinear passive plant. We replace the linear spring between the carts with a piece-wise linear one, as shown in Figure 6. The force-displacement characteristic has slope $1$ for small displacements ($|\Delta x|\leq 0.5$) and slope $2$ for large displacements ($|\Delta x|\geq 0.5$).
Figure 6: Nonlinear spring.
To deal with the nonlinearity, we also reduce the aggressiveness of the reference model, by taking $T_{1}=1$. The iFIR controller is designed using (2) and (4), based on data from the nonlinear plant. Performance degrades but stability is guaranteed from passivity. The iFIR design still outperforms the optimal PID controller (based on nonlinear plant data).
V Conclusions
We have discussed a new data-driven approach for passive iFIR controllers. Their design is based on convex optimization, combining virtual reference feedback tuning with passivity constraints. Our approach does not require any plant model nor large datasets, and the plant can be nonlinear. For passive plants, stability of the closed loop is guaranteed by construction. Future work will focus on MIMO systems and nonlinear controllers.
@@ V Conclusions
Figure 7: Bode diagrams. Red: target reference model. Black: iFIR-based (2),(4) closed-loop transfer function from $r$ to $y$ Blue-continuous: PID based closed-loop transfer function. Blue-dashed: plant transfer function.
@@
Figure 8: Step responses of the controlled nonlinear plant. Red: target response. Black: nonlinear closed-loop response with iFIR controller. Blue-continuous: nonlinear closed-loop response with PID. Blue-dashed: nonlinear open loop.

### v2

@@ 4.1 Fitting of passive FIR filters
where $C_{1}$ is passive, and $C_{2}$ and $C_{3}$ show increase shortage of passivity. For each target, we take $\gamma=0$ and (5) paired with passivity constraints. The input/output data $(e,u)$ are generated as $u=C_{i}e$ for $i\in\{1,2,3\}$. The probing signal is a filtered step given by $e=\mathcal{L}^{-1}\left(\frac{1}{0.2s+1}\frac{1}{s}\right)$, where $\mathcal{L}^{-1}$ is the inverse Laplace transform. Results are shown in Figure 2. The top-left plot shows that all three approaches give perfect fitting. The top-right plot shows the results for conservative $\epsilon$ (10d). Indeed, this parameter can be used to enforce an excess of passivity, to compensate for shortages in the plant. The left and right bottom plots are related to the non-passive targets $C_{2}$ and $C_{3}$. All three approaches achieve a passive FIR filter. As expected, the match with the target systems degrades as the passivity shortage increases. These results are obtained using CVXPY [13] with MOSEK solver [24].
4.2 Velocity control of a compliant mechanical system
@@
Consider the two-cart plant of Figure 3, modeling a truck pulling a heavy load. The cart masses are $m_{1}=3.0\mbox{ kg},m_{2}=0.5\mbox{ kg}$. $k_{12}=1.0$ Ns/m is the spring stiffness and $c_{12}=1.05$ Ns/m is the damping coefficient. Each cart is also affected by additional dissipation of the form $c\dot{x}_{i}$, where $c=0.5$ Ns/m and $\dot{x}_{i}$ is the absolute velocity of the cart $i\in\{1,2\}$. The block diagram for this closed-loop system is represented in Figure 1. $u$ is the applied force and $y$ is the velocity output.
@@
We consider the two reference models $M_{r1}$ and $M_{r2}$
${\color[rgb]{0,0,0}M_{ri}(s)}$ ${\color[rgb]{0,0,0}=\frac{1+0.1s}{1+2\zeta sT_{i}+s^{2}T_{i}^{2}}\qquad i=1,2}$
where $\zeta=1$, $T_{1}=0.25$s, $T_{2}=0.125$s. $M_{r2}$ is more aggressive because it has a higher cutoff frequency. The plant and the reference models are discretized with the zero-order-hold method at a sampling time $T_{s}$ of $0.05$s. The plant is excited in open-loop with $u(t)=\sum_{i=1}^{10}\sin(\omega_{i}T_{s}t)$ where $\omega_{i}$ linearly spans the frequency range of $[0.5,10]$ rad/s, and $t\in\{0,1,...,2000\}$. We assume that there is no noise in the input and output of the plant. Standard techniques such as instrumental variables [10, Section 4] can be used in the noisy case. The data $(e,u)$ for the controller are obtained as discussed in Section 2 and filtered with the filters $M_{r1}(z)$ and $M_{r2}(z)$.
For $M_{r1}$, the computation times for several iFIR controllers of order $m\in\{50,150,250,350\}$ are summarized in Table 1. The times refer to the solution of the optimization problem on a standard Apple Silicon M2. Compilation times are excluded. For $m=350$, the KYP approach takes more than one hour. This shows the advantage of (10) and (14) over (7).
@@
Thm. 3.1 0.24 s 54.03 s 705.44 s ///
Thm. 3.5, $n=m$ 0.07 s 0.51 s 2.81 s 4.88 s
Thm. 3.5, $n=1.5m$ 0.12 s 1.88 s 5.21 s 12.18 s
Thm. 3.5, $n=2m$ 0.17 s 3.36 s 10.62 s 61.14 s
Thm. 3.7, $M=m$ 0.05 s 0.14 s 0.17 s 0.25 s
Thm. 3.7, $M=1.5m$ 0.04 s 0.14 s 0.17 s 0.24 s
Thm. 3.7, $M=2m$ 0.05 s 0.14 s 0.20 s 0.26 s
Table 1: The computation times of passive iFIR design.
@@
The Nyquist plots of the FIR part of the iFIR controllers are shown in Figure 4. Since plots are similar, only the Toeplitz case is shown for reasons of space. We have taken $m=350$, $n=2m$ for (10) and $M=2m$ for (14). $\epsilon$ is chosen heuristically. On the right-part of the complex plane, both passive iFIR controllers (black) show excellent matching with the optimal unconstrained ones (blue) .
@@
Figure 4: Nyquist plots of the FIR part of the iFIR controller. Blue - unconstrained (2); black - constrained (2), (10). Left: $M_{r1}$; Right: $M_{r2}$.
The iFIR controllers are tested in simulation, as shown in Figure 5. The figure shows a comparison between passive iFIR controllers and the optimal PID controller $C(z)=K_{p}+K_{d}\frac{T_{s}}{1-z^{-1}}+K_{i}\frac{z-1}{zT_{s}}$ where $[K_{p},K_{d},K_{i}]$ correspond to $[0.8051,4.4090,0.0068]$ and $[2.6142,10.2330,0.0232]$ for $M_{r1}$ and $M_{r2}$, respectively. The PID gains are constrained to be non-negative (for passivity) and tuned using VRFT. For space reasons, we do not develop a comparison with non-passive iFIR controllers and PID controllers.
@@
Figure 5: Step responses. Red: target response. Black: closed-loop response with passive iFIR controller (2), (10). Blue-continuous: closed-loop response with PID. Blue-dashed: open loop. Left: $M_{r1}$; Right: $M_{r2}$.
The closed-loop response of the two iFIR controllers is superior, with an almost exact matching to the target model. This is further illustrated by the Bode diagrams of the closed-loop transfer functions, in Figure 6.
@@
Figure 6: Bode diagrams when $M_{r1}$ is used. Left: magnitude plot. Right: phase plot. Red: target reference model. Black: iFIR-based (2), (10) closed loop. Blue-continuous: PID based closed loop. Blue-dotted: plant. Similar results hold for $M_{r2}$.
We also design and test the iFIR controller on a nonlinear passive plant. We replace the linear spring between the carts with a piece-wise linear one. The force-displacement characteristic has slope $1$ for small displacements ($|\Delta x|\leq 0.5$) and $2$ for large displacements ($|\Delta x|\geq 0.5$). To deal with the nonlinearity, we reduce the aggressiveness of the reference model, by taking the reference model to be $\frac{1+0.1s}{1+2\zeta sT+s^{2}T^{2}},\zeta=1$, $T=1$s. The iFIR controller is designed using (2) and (14), based on nonlinear plant data. Performance degrades but stability is guaranteed. The iFIR design outperforms the optimal PID, as shown in Figure 7.
@@
Figure 7: Left: step responses of the controlled nonlinear plant. Red: target response. Black: nonlinear closed-loop response with iFIR controller. Blue-continuous: nonlinear closed-loop response with PID. Blue-dashed: nonlinear open loop. Right: nonlinear spring force-displacement characteristics.
5 Conclusions
We have discussed a new data-driven approach for passive iFIR controllers. Their design is based on convex optimization, combining virtual reference feedback tuning with passivity constraints. Our approach does not require any plant model or large datasets, and the plant can be nonlinear. For passive plants, the stability of the closed loop is guaranteed by construction. Future work will extend the design to MIMO systems and to nonlinear controllers.